OpenStudy (anonymous):

for the equation -4y=8x, what is the constant of variation Math, help??

OpenStudy (immanuelv):

for example y = 3x + 2 2 is the constant in this case there is no constant

OpenStudy (anonymous):

I have the option of a) -2 b) -4 c) 1 D) 2

OpenStudy (immanuelv):

are you sure you wrote your equation right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Heres another one you can maybe help me with Suppose y varies directly with x, and y= 24 when x=8 . What is the value of y when x=10? a) 10 b)30 c)80 d)240

OpenStudy (immanuelv):

3y= x so if x = 8, y = 8*3 then if x = 10 then y= 10*3